# SAVeD: Semantically Aware Version Discovery

### v1

@@ 1. Introduction
In any field revolving around data, it is reasonable to assume that outside of a heavily controlled environment, data is ”unclean” and ”unorganized”, that is, it lacks structure. In both corporate and academic settings, datasets often require some transformation in order to be of use to an engineer or developer. However, the transformations performed on the same piece of data vary wildly, even on a set of data considered to be ”solved”. (CaiZhu2015)
@@
Discovering these versions is crucial for avoiding redundant work and leveraging existing data science efforts. The discovery of these valid tables is called Version Discovery. This problem represents the first explicit exploration of automated version discovery (10.14778/3583140.3583169), addressing a gap in prior work on data versioning and management.
Version discovery can be viewed as a dataset discovery task. It exists within the broader landscape of dataset discovery methods such as keyword search (4430467) and query search (DBLP:journals/corr/GaoC17). More recently, dataset discovery has focused on the tasks of Table Union Search (10.14778/3192965.3192973) and Joinable table search (10.1145/3299869.3300065), both of which focus on the interaction between tables in a data lake. In the realm of version discovery, prior work has focused on creating benchmarks of table versions using seed tables (10.14778/3583140.3583169) and creating benchmarks using Large Language Models (LLMs) (10.1145/3627673.3679157). However, no work has been performed (to the authors’ knowledge) on version discovery itself, that is, finding whether existing modifications have been made to a given dataset already. For the task of table union search, contrastive learning (fan2023semanticsawaredatasetdiscoverydata) has been used as a means of augmenting the search in an efficient manner. We posit that contrastive learning on table pairs can also be used as a means of augmentation to discover realistic, internal-driven transformations for tables in a data lake, creating a framework for semantic-aware version discovery.
@@ 2.1. Data Versioning
Data versioning has been traditionally approached from a metadata-driven perspective. The foundational work of Sciore (Sciore94) established theoretical frameworks for database versioning, while Bhardwaj, et. al’s work on DataHub (bhardwaj2014datahubcollaborativedatascience) identified key challenges in collaborative data science environments. Explain Da V(10.14778/3583140.3583169) introduced the concept of semantic versioning in databases, establishing benchmarks for version identification but assuming prior knowledge of version relationships.
Modern version control systems like DVC (ruslan_kuprieiev_2025_17037377), Pachyderm (Pachyderm), and even AutoML plaforms like MLflow(Zaharia2018AcceleratingTM) provide infrastructure for tracking data changes but require explicit version declarations and metadata maintenance. Git (git) extends traditional version control to data but struggles with large binary files, relying on approaches like its Large File System (LFS). Recent approaches like Huggingface’s Xet have bypassed this limitation, but inherently lack semantic understanding.
The database community has explored temporal databases (404027) and schema evolution (inproceedingsmanousis) extensively, but these approaches focus on structural changes rather than semantic relationships. Work by Fox et al.(10.1145/3627673.3679157) explored LLM-based approaches for generating synthetic table versions using prompt engineering and GPT models, demonstrating the complexity of realistic version creation but focusing on benchmark generation rather than discovery. Shraga and Miller (10.14778/3583140.3583169) developed Explain-Da-V and established the Semantic Versioning in Databases (SDVB) benchmark for explaining dataset changes but assumed known version relationships.
@@ 2.2. Data Discovery
Table discovery and search have been extensively studied in database research, evolving from early schema matching approaches to modern semantic-aware systems. Traditional approaches like COMA (inproceedings) and Similarity Flooding (10.5555/876875.879024) focused on structural and linguistic similarity.
@@
Web table discovery systems emerged with the growth of HTML tables. WebTables (10.14778/1453856.1453916) and subsequent work by Lehmberg et al.(10.1145/2872518.2889386) demonstrated large-scale table extraction and matching from web sources. However, these approaches primarily address structural heterogeneity rather than semantic versioning relationships.
@@
Recent advances in deep learning leverage deep semantic relationships for table understanding. TaBERT(yin2020tabertpretrainingjointunderstanding) and TAPAS(Herzig_2020) learn joint representations of tables and natural language, enabling question-answering over tabular data. Starmie(fan2023semanticsawaredatasetdiscoverydata) uses pre-trained RoBERTa models for table classification and retrieval tasks, achieving state-of-the-art performance on benchmark datasets like the aforementioned Semantic Versioning in Databases dataset.
@@
The advent of foundation models has spawned new approaches: TAPEX (liu2022tapextablepretraininglearning) learns to execute SQL queries by treating tables as structured text and the ToTTo (parikh2020tottocontrolledtabletotextgeneration) benchmark addresses table-to-text generation. While these approaches excel at table comprehension tasks, they require structured supervision and operate fundamentally differently from version discovery, which needs to learn semantic relationships from unlabeled table pairs without assuming well-defined task objectives.
@@
The subtask of joinability detection represents a related but distinct problem. JOSIE (10.1145/3299869.3300065) and LSH Forest (10.1145/1060745.1060840) focus on finding tables that can be joined based on overlapping values, while our work addresses semantic equivalence despite structural differences. Starmie (fan2023semanticsawaredatasetdiscoverydata) represents the most closely related work to our approach. It uses pre-trained RoBERTa models with contrastive learning for table understanding tasks, achieving state-of-the-art results on various benchmarks including the Semantic Versioning in Databases dataset. However, Starmie focuses on general table discovery and requires supervised fine-tuning for specific tasks. In contrast, SAVeD is specifically designed for version discovery through unsupervised contrastive learning, with custom augmentations tailored to capture version relationships without labeled data. For a detailed discussion of contrastive learning methods in data integration and discovery, see Section 4.1. (fan2023semanticsawaredatasetdiscoverydata) has used contrastive learning along with labeled pairs to learn a semantic-aware representation of tables for the task of joinability. While the methodology is solid, Starmie’s contrastive objective is anchored in supervised pairs curated for joinability and table discovery, As a result, its embedding space is shaped to capture cross-table semantic similarity rather than intra-dataset evolutionary changes. Second, the augmentations used in Starmie are designed to enhance robustness to surface-level noise (e.g., column permutations or name variations), but they do not simulate version-like transformations such as incremental schema evolution, value drift, or attribute-level refinements. These differences cause Starmie’s embeddings to collapse version-specific nuances that are essential for version discovery. By contrast, SAVeD’s unsupervised pipeline forces the model to learn the kinds of fine-grained relationships that supervised table-discovery pre-training tends to overlook.
@@ 3.2. Approach
This defines a tolerance relation (zeeman1962topology) on $\mathcal{T}$, which is weaker than an equivalence relation due to the approximate nature of $\theta$.
@@ 3.3. Motivating Example
This example demonstrates the core challenge of version discovery: automatic methods must be invariant to superficial structural changes (encoding, ordering, column addition/removal) while sensitive to semantic changes that alter the meaning of the data. The goal of version discovery is to provide an unsupervised solution to this issue, and as opposed to related table search (10.1145/2213836.2213962), version semantics need to be explicitly captured. All prior exploration in Data Versioning has proceeded under the assumption that the relationship between tables is known beforehand, but this must be discarded beforehand in order to best learn a relationship between $T$ and $T^{\prime}$, taking into account the lack of labeled examples, and the motivations behind the problem.
@@ 4.1. Preliminaries: Contrastive Learning
Contrastive learning has emerged as a fundamental paradigm for self-supervised representation learning, with theoretical foundations rooted in metric learning (1640964) and noise contrastive estimation (JMLR:v13:gutmann12a). The core principle involves learning representations that pull semantically similar samples together while pushing dissimilar samples apart in the embedding space.
@@
Early work in computer vision established the effectiveness of contrastive approaches. Deep Metric Learning (sym11091066) and Triplet Loss(Schroff2015FaceNetAU) demonstrated that carefully constructed positive and negative pairs could learn discriminative embeddings. The breakthrough came with SimCLR(chen2020simpleframeworkcontrastivelearning), which showed that simple data augmentations combined with large batch sizes and strong data augmentation on the fly could achieve state-of-the-art self-supervised performance, rivaling supervised methods on ImageNet.
@@
Subsequent developments expanded contrastive learning across domains. In natural language processing, SimCSE (gao2022simcsesimplecontrastivelearning) applied contrastive learning to sentence embeddings, while ConSERT(yan2021consertcontrastiveframeworkselfsupervised) and DiffCSE(chuang2022diffcsedifferencebasedcontrastivelearning) refined the approach with improved augmentation strategies. The theoretical understanding of contrastive learning has evolved significantly. Wang and Isola(wang2022understandingcontrastiverepresentationlearning) provided theoretical analysis showing that contrastive learning implicitly performs spectral clustering. HaoChen et al. (DBLP:journals/corr/abs-2106-04156) established generalization bounds, while Tosh et al. (tosh2021contrastivelearningmultiviewredundancy) analyzed the optimization landscape of contrastive objectives.
@@
Recent advances have focused on addressing key limitations: MoCo(he2020momentumcontrastunsupervisedvisual) introduced momentum-based negative sampling to enable larger effective batch sizes, SwAV(caron2021unsupervisedlearningvisualfeatures) eliminated the need for negative pairs through clustering, and BYOL(grill2020bootstraplatentnewapproach) achieved contrastive learning without explicit negative samples.
@@
In tabular data, applications have been limited. VIME(NEURIPS2020_7d97667a) applied contrastive pretraining to tabular data for downstream prediction tasks, while TabNet(arik2020tabnetattentiveinterpretabletabular) incorporated attention mechanisms but didn’t leverage contrastive learning. More recent work by Bahri et al.(bahri2022scarfselfsupervisedcontrastivelearning) introduced SCARF for tabular representation learning, but focused on supervised fine-tuning rather than discovery tasks. Most recently, Starmie (fan2023semanticsawaredatasetdiscoverydata) leveraged contrastive learning for dataset discovery. Our approach to version discovery centers around using contrastive learning techniques, which allow models a label-free way to distinguish between similar and dissimilar data points. We implement a SimCLR-style framework (chen2020simpleframeworkcontrastivelearning), generating positive and negative pairs from a given batch of samples, and is created in a way to address the challenges of locating semantically similar table versions (”versions” of a dataset). Since version discovery requires identifying different representations of the same data via transformation, supervised methods for the task of version discovery are impractical, due to the lack of labeled data for the task.
@@ 4.4. NT-Xent Contrastive Loss
where $\tau=0.7$ is our temperature parameter, as per Fan, et. al (fan2023semanticsawaredatasetdiscoverydata)
@@
This formulation ensures that positive pairs (augmented views of the same table) are pulled together in the embedding space while being pushed away from all other tables in the batch, enabling effective version relationship learning. This also ensures that perturbations which can occur naturally (such as those we established earlier) can be learned, as they are closed in embedding space than those that cannot be. Following the analysis of Wang and Isola (wang2022understandingcontrastiverepresentationlearning), our contrastive objective can be inferred to perform a form of spectral clustering on the table similarity graph. Let $G=(V,E,W)$ be a weighted graph where $V$ represents tables, $E$ contains edges between semantically similar tables, and $W$ contains similarity weights.
@@
Under mild conditions, this objective converges to representations that preserve the dominant eigenvectors of the normalized graph Laplacian $L=I-D^{-1/2}WD^{-1/2}$, where $D$ is the degree matrix(868688). This convergence implicitly preserves semantics via spectral clustering, providing a basis for our contrastive method that can serve to provide a powerful embedding space even with minimal amounts of labeled data.
@@ 4.5. SAVeD Architecture
Transformer Encoder: We implement a multi-layer transformer encoder with $N$ layers, each containing multi-head self-attention with $h$ heads and position-wise feedforward networks. Following Vaswani et al. (DBLP:journals/corr/VaswaniSPUJGKP17), we include residual connections and layer normalization.
@@ 5.3. Training
We trained with the AdamW optimizer (Loshchilov2017DecoupledWD) with the a learning rate of $2.3\times 10^{-4}$ (optimized via hyperparameter search), weight decay of $5.7\times 10^{-5}$ for L2 regularization, beta parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$, and a batch size of 32, which is chosen for a balance of memory efficiency and gradient stability). Our training protocol involves a linear learning rate warm-up over first 10% of training steps, followed by the main contrastive learning process. To mitigate overfitting, we add patience-based early stopping which monitors changes in the validation loss. Our optimal model is selected on the minimization of validation loss and maximization of separation. For cross-validation, we implement a stratified splitting approach, where 70% of tables per dataset are used for training, 15% of the tables per dataset are used for for hyperparameter optimization and early stopping, and the final 15% are used for the final evaluation and are completely held out during training. Each training batch contains tables from multiple datasets to encourage generalization. Positive pairs are constructed within-batch through augmentation, while negative pairs are formed from all other tables in the batch. We also implement gradient clipping (max norm = 1.0) to prevent exploding gradients during contrastive learning, particularly important given the temperature scaling in our loss function.

### v2

@@ 1. Introduction
In any field revolving around data, it is reasonable to assume that outside of a heavily controlled environment, data is ”unclean” and ”unorganized”, that is, it lacks structure. In both corporate and academic settings, datasets often require some transformation in order to be of use to an engineer or developer. However, the transformations performed on the same piece of data vary wildly, even on a set of data considered to be ”solved”. (Cai and Zhu, 2015)
@@
Discovering these versions is crucial for avoiding redundant work and leveraging existing data science efforts. The discovery of these valid tables is called Version Discovery. This problem represents the first explicit exploration of automated version discovery (Shraga and Miller, 2023), addressing a gap in prior work on data versioning and management.
Version discovery can be viewed as a dataset discovery task. It exists within the broader landscape of dataset discovery methods such as keyword search (Gupta, 2007) and query search (Gao and Callan, 2017). More recently, dataset discovery has focused on the tasks of Table Union Search (Nargesian et al., 2018) and Joinable table search (Zhu et al., 2019), both of which focus on the interaction between tables in a data lake. In the realm of version discovery, prior work has focused on creating benchmarks of table versions using seed tables (Shraga and Miller, 2023) and creating benchmarks using Large Language Models (LLMs) (Fox et al., 2024). However, no work has been performed (to the authors’ knowledge) on version discovery itself, that is, finding whether existing modifications have been made to a given dataset already. For the task of table union search, contrastive learning (Fan et al., 2023) has been used as a means of augmenting the search in an efficient manner. We posit that contrastive learning on table pairs can also be used as a means of augmentation to discover realistic, internal-driven transformations for tables in a data lake, creating a framework for semantic-aware version discovery.
@@ 2.1. Data Versioning
Data versioning has been traditionally approached from a metadata-driven perspective. The foundational work of Sciore (Sciore, 1994) established theoretical frameworks for database versioning, while Bhardwaj, et. al’s work on DataHub (Bhardwaj et al., 2014) identified key challenges in collaborative data science environments. Explain Da V(Shraga and Miller, 2023) introduced the concept of semantic versioning in databases, establishing benchmarks for version identification but assuming prior knowledge of version relationships.
Modern version control systems like DVC (Kuprieiev et al., 2025), Pachyderm (Pachyderm. Inc., 2025), and even AutoML plaforms like MLflow(Zaharia et al., 2018) provide infrastructure for tracking data changes but require explicit version declarations and metadata maintenance. Git (Torvalds and Hamano, 2005) extends traditional version control to data but struggles with large binary files, relying on approaches like its Large File System (LFS). Recent approaches like Huggingface’s Xet have bypassed this limitation, but inherently lack semantic understanding.
The database community has explored temporal databases (Ozsoyoglu and Snodgrass, 1995) and schema evolution (Manousis et al., 2016) extensively, but these approaches focus on structural changes rather than semantic relationships. Work by Fox et al.(Fox et al., 2024) explored LLM-based approaches for generating synthetic table versions using prompt engineering and GPT models, demonstrating the complexity of realistic version creation but focusing on benchmark generation rather than discovery. Shraga and Miller (Shraga and Miller, 2023) developed Explain-Da-V and established the Semantic Versioning in Databases Benchmark (SDVB) for explaining dataset changes but assumed known version relationships.
@@ 2.2. Data Discovery
Table discovery and search have been extensively studied in database research, evolving from early schema matching approaches to modern semantic-aware systems. Traditional approaches like COMA (Aumueller et al., 2005) and Similarity Flooding (Melnik et al., 2002) focused on structural and linguistic similarity.
@@
Web table discovery systems emerged with the growth of HTML tables. WebTables (Cafarella et al., 2008) and subsequent work by Lehmberg et al.(Lehmberg et al., 2016) demonstrated large-scale table extraction and matching from web sources. However, these approaches primarily address structural heterogeneity rather than semantic versioning relationships.
@@
Recent advances in deep learning leverage deep semantic relationships for table understanding. TaBERT(Yin et al., 2020) and TAPAS(Herzig et al., 2020) learn joint representations of tables and natural language, enabling question-answering over tabular data. Starmie(Fan et al., 2023) uses pre-trained RoBERTa models for table classification and retrieval tasks, achieving state-of-the-art performance on benchmark datasets like the aforementioned Semantic Versioning in Databases dataset.
@@
The advent of foundation models has spawned new approaches: TAPEX (Liu et al., 2022) learns to execute SQL queries by treating tables as structured text and the ToTTo (Parikh et al., 2020) benchmark addresses table-to-text generation. While these approaches excel at table comprehension tasks, they require structured supervision and operate fundamentally differently from version discovery, which needs to learn semantic relationships from unlabeled table pairs without assuming well-defined task objectives.
@@
The subtask of joinability detection represents a related but distinct problem. JOSIE (Zhu et al., 2019) and LSH Forest (Bawa et al., 2005) focus on finding tables that can be joined based on overlapping values, while our work addresses semantic equivalence despite structural differences. Starmie (Fan et al., 2023) represents the most closely related work to our approach. It uses pre-trained RoBERTa models with contrastive learning for table understanding tasks, achieving state-of-the-art results on various benchmarks including the Semantic Versioning in Databases dataset. However, Starmie focuses on general table discovery and requires supervised fine-tuning for specific tasks. In contrast, SAVeD is specifically designed for version discovery through unsupervised contrastive learning, with custom augmentations tailored to capture version relationships without labeled data. For a detailed discussion of contrastive learning methods in data integration and discovery, see Section 4.1. (Fan et al., 2023) has used contrastive learning along with labeled pairs to learn a semantic-aware representation of tables for the task of joinability. While the methodology is solid, Starmie’s contrastive objective is anchored in supervised pairs curated for joinability and table discovery, As a result, its embedding space is shaped to capture cross-table semantic similarity rather than intra-dataset evolutionary changes. Second, the augmentations used in Starmie are designed to enhance robustness to surface-level noise (e.g., column permutations or name variations), but they do not simulate version-like transformations such as incremental schema evolution, value drift, or attribute-level refinements. These differences cause Starmie’s embeddings to collapse version-specific nuances that are essential for version discovery. By contrast, SAVeD’s unsupervised pipeline forces the model to learn the kinds of fine-grained relationships that supervised table-discovery pre-training tends to overlook.
@@ 3.2. Approach
This defines a tolerance relation (Zeeman, 1962) on $\mathcal{T}$, which is weaker than an equivalence relation due to the approximate nature of $\theta$.
@@ 3.3. Motivating Example
This example demonstrates the core challenge of version discovery: automatic methods must be invariant to superficial structural changes (encoding, ordering, column addition/removal) while sensitive to semantic changes that alter the meaning of the data. The goal of version discovery is to provide an unsupervised solution to this issue, and as opposed to related table search (Das Sarma et al., 2012), version semantics need to be explicitly captured. All prior exploration in Data Versioning has proceeded under the assumption that the relationship between tables is known beforehand, but this must be discarded beforehand in order to best learn a relationship between $T$ and $T^{\prime}$, taking into account the lack of labeled examples, and the motivations behind the problem.
@@ 4.1. Preliminaries: Contrastive Learning
Contrastive learning has emerged as a fundamental paradigm for self-supervised representation learning, with theoretical foundations rooted in metric learning (Hadsell et al., 2006) and noise contrastive estimation (Gutmann and Hyvärinen, 2012). The core principle involves learning representations that pull semantically similar samples together while pushing dissimilar samples apart in the embedding space.
@@
Early work in computer vision established the effectiveness of contrastive approaches. Deep Metric Learning (KAYA and BİLGE, 2019) and Triplet Loss(Schroff et al., 2015) demonstrated that carefully constructed positive and negative pairs could learn discriminative embeddings. The breakthrough came with SimCLR(Chen et al., 2020), which showed that simple data augmentations combined with large batch sizes and strong data augmentation on the fly could achieve state-of-the-art self-supervised performance, rivaling supervised methods on ImageNet.
@@
Subsequent developments expanded contrastive learning across domains. In natural language processing, SimCSE (Gao et al., 2022) applied contrastive learning to sentence embeddings, while ConSERT(Yan et al., 2021) and DiffCSE(Chuang et al., 2022) refined the approach with improved augmentation strategies. The theoretical understanding of contrastive learning has evolved significantly. Wang and Isola(Wang and Isola, 2022) provided theoretical analysis showing that contrastive learning implicitly performs spectral clustering. HaoChen et al. (HaoChen et al., 2021) established generalization bounds, while Tosh et al. (Tosh et al., 2021) analyzed the optimization landscape of contrastive objectives.
@@
Recent advances have focused on addressing key limitations: MoCo(He et al., 2020) introduced momentum-based negative sampling to enable larger effective batch sizes, SwAV(Caron et al., 2021) eliminated the need for negative pairs through clustering, and BYOL(Grill et al., 2020) achieved contrastive learning without explicit negative samples.
@@
In tabular data, applications have been limited. VIME(Yoon et al., 2020) applied contrastive pretraining to tabular data for downstream prediction tasks, while TabNet(Arik and Pfister, 2020) incorporated attention mechanisms but didn’t leverage contrastive learning. More recent work by Bahri et al.(Bahri et al., 2022) introduced SCARF for tabular representation learning, but focused on supervised fine-tuning rather than discovery tasks. Most recently, Starmie (Fan et al., 2023) leveraged contrastive learning for dataset discovery. Our approach to version discovery centers around using contrastive learning techniques, which allow models a label-free way to distinguish between similar and dissimilar data points. We implement a SimCLR-style framework (Chen et al., 2020), generating positive and negative pairs from a given batch of samples, and is created in a way to address the challenges of locating semantically similar table versions (”versions” of a dataset). Since version discovery requires identifying different representations of the same data via transformation, supervised methods for the task of version discovery are impractical, due to the lack of labeled data for the task.
@@ 4.4. NT-Xent Contrastive Loss
where $\tau=0.7$ is our temperature parameter, as per Fan, et. al (Fan et al., 2023)
@@
This formulation ensures that positive pairs (augmented views of the same table) are pulled together in the embedding space while being pushed away from all other tables in the batch, enabling effective version relationship learning. This also ensures that perturbations which can occur naturally (such as those we established earlier) can be learned, as they are closed in embedding space than those that cannot be. Following the analysis of Wang and Isola (Wang and Isola, 2022), our contrastive objective can be inferred to perform a form of spectral clustering on the table similarity graph. Let $G=(V,E,W)$ be a weighted graph where $V$ represents tables, $E$ contains edges between semantically similar tables, and $W$ contains similarity weights.
@@
Under mild conditions, this objective converges to representations that preserve the dominant eigenvectors of the normalized graph Laplacian $L=I-D^{-1/2}WD^{-1/2}$, where $D$ is the degree matrix(Shi and Malik, 2000). This convergence implicitly preserves semantics via spectral clustering, providing a basis for our contrastive method that can serve to provide a powerful embedding space even with minimal amounts of labeled data.
@@ 4.5. SAVeD Architecture
Transformer Encoder: We implement a multi-layer transformer encoder with $N$ layers, each containing multi-head self-attention with $h$ heads and position-wise feedforward networks. Following Vaswani et al. (Vaswani et al., 2017), we include residual connections and layer normalization.
@@ 5.3. Training
We trained with the AdamW optimizer (Loshchilov and Hutter, 2017) with the a learning rate of $2.3\times 10^{-4}$ (optimized via hyperparameter search), weight decay of $5.7\times 10^{-5}$ for L2 regularization, beta parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$, and a batch size of 32, which is chosen for a balance of memory efficiency and gradient stability). Our training protocol involves a linear learning rate warm-up over first 10% of training steps, followed by the main contrastive learning process. To mitigate overfitting, we add patience-based early stopping which monitors changes in the validation loss. Our optimal model is selected on the minimization of validation loss and maximization of separation. For cross-validation, we implement a stratified splitting approach, where 70% of tables per dataset are used for training, 15% of the tables per dataset are used for for hyperparameter optimization and early stopping, and the final 15% are used for the final evaluation and are completely held out during training. Each training batch contains tables from multiple datasets to encourage generalization. Positive pairs are constructed within-batch through augmentation, while negative pairs are formed from all other tables in the batch. We also implement gradient clipping (max norm = 1.0) to prevent exploding gradients during contrastive learning, particularly important given the temperature scaling in our loss function.